# Quasi-symmetry Constrained Spin Ferromagnetism in Altermagnets

### v1

@@ Abstract
Altermagnets break time-reversal symmetry and their spin-orbit coupling (SOC) allow for an anomalous Hall effect (AHE) that depends on the direction of the Néel ordering vector. The AHE and the ferromagnetic spin moment share the same symmetry and hence are usually proportional. However, density functional theory (DFT) calculations find that the AHE exists with negligible ferromagnetic spin moment for some compounds, whereas it reaches sizable values for other altermagnets. By examining realistic minimal models for altermagnetism in which the DFT phenomenology is captured, we uncover a general SOC-enabled quasi-symmetry that provides a natural explanation for the amplitude of the ferromagnetic spin moment across the vast range of different altermagnetic materials. Additionally, we derive analytic expressions for the magnetic anisotropy energy, providing a simple means to identify the preferred Néel vector orientation for altermagnets.
@@
Introduction.– The Hall effect has been a frontier theme in condensed matter physics for more than half a century. Fundamental understanding of the quantized Hall conductance and the anomalous Hall effect (AHE) has helped pave the way for topological classifications of quantum matter and the importance of Berry curvature in transport properties Nagaosa et al. (2010). More recently, this has been generalized to nonlinear transport, which has been described in terms of multipoles of the quantum geometric tensor Sodemann and Fu (2015); Anyuan Gao, Yu-Fei Liu, Jian-Xiang Qiu, Barun Ghosh, Thaís V. Trevisan, Yugo Onishi, Chaowei Hu, Tiema Qian, Hung-Ju Tien, Shao-Wen Chen, Mengqi Huang, Damien Bérubé, Houchen Li, Christian Tzschaschel, Thao Dinh, et al. (2023); Kaplan et al. (2024); Fang et al. (2024); Du et al. (2021); Wang et al. (2023). The discovery of altermagnetism (AM) and its associated large antiferromagnetic AHE is yet another testimony to this development Lee et al. (2024); Krempaský et al. (2024); Osumi et al. (2024); Li et al. (2024a); Yang et al. (2024); Reimers et al. (2024); Ding et al. (2024); Feng et al. (2022); Kluczyk et al. (2024); Gonzalez Betancourt et al. (2023), which may open new opportunities for devices utilizing dissipationless transport and spintronics technologies Šmejkal et al. (2020, 2022a, 2022b, 2022c); Fang et al. (2024).
@@
The AHE in AM is only nonzero for certain directions of the Néel vector and in the presence of relativistic symmetry-lowering spin-orbit coupling (SOC). This, in addition, induces weak ferromagnetism (FM), which satisfies the same symmetry properties as the AHE McClarty and Rau (2024); Xiao et al. (2024). This implies that, in principle, FM and AHE share the same dependence on the orientation of the Néel vector Šmejkal et al. (2022c); Fernandes et al. (2024); Cheong and Huang (2024). However, from DFT it is found that the relative amplitude of the AHE transport coefficient and the FM spin moment depends strongly on the particular AM material under investigation. For example, for the materials RuO2 and MnTe, DFT finds large AHE of 40 S/cm and 300 S/cm, respectively, while the possible associated FM spin moment is either vanishing for RuO2 and very small of order $10^{-4}\mu_{B}$ per Mn for MnTe Šmejkal et al. (2020); Mazin (2024); Gonzalez Betancourt et al. (2023). In the case of MnTe and RuO2, weak FM was also reported experimentally Berlijn et al. (2017); Kluczyk et al. (2024), though the AM ground state of the latter material is still debated Hiraishi et al. (2024); Keßler et al. (2024). A weak FM spin moment has also been identified by DFT in CrSb Autieri et al. (2023). By contrast, in other AM, such as FeSb2, DFT finds a large AHE of 143 S/cm and a relatively larger FM moment Mazin et al. (2021); Attias et al. (2024). Similarly, the material RuF4 has also been predicted by DFT to exhibit a large FM spin moment of 0.22 $\mu_{B}$ on the Ru sites Milivojević et al. (2024).
@@
The results highlighted above reveal that the interplay between AHE and weak FM in AM materials is not well understood. To gain deeper insight into the weak FM properties of AM it is crucial to analyze the role of SOC using realistic models. Here, by examining realistic microscopic models for AM in which the DFT phenomenology is captured Roig et al. (2024), we identify which AM states have a large SOC induced AHE but a small FM spin moment. In addition, we find that our model-based results are generally valid. Specifically, they apply to all microscopic theories in which AM appears as a consequence of the exchange interaction. To show this, we identify a general SOC-enabled quasi-symmetry Li et al. (2024b); Guo et al. (2022) that establishes for which AM symmetries the symmetry-allowed FM is large or vanishingly small. Finally, the presence of SOC in AM breaks the spin-space degeneracy and leads to a preferred direction for the Néel vector. For example, DFT calculations have revealed that the moments are out-of-plane in RuO2 Feng et al. (2022), while in MnTe, FeSb2, Nb2FeB2 and Ta2FeB2 the moments are predicted to be in-plane Mazin et al. (2021); Mazin (2024); Hou et al. (2023). By examining analytic expressions for the Landau coefficients derived from realistic microscopic models, we elucidate how the structure of SOC allows for the AM anisotropy energy to be understood.
@@
Interplay between magnetization and Néel order.– Introducing ${\vec{N}}$ as the Néel order and ${\vec{M}}$ as the magnetization, the general form of the free energy density in the presence of SOC can be written as
@@
where $a_{N}$, $b_{N}$, $a_{M}$ and $c_{ij}$ are temperature-dependent Landau coefficients, $\vec{h}$ is an applied field, and $s_{1}$ and $s_{2}$ are the coefficients determining the magnetic anisotropy energy due to SOC. Note that the bilinear coupling between the two orders $M_{i}$ and $N_{j}$ is only allowed in the presence of SOC McClarty and Rau (2024). This free energy describes orthorhombic or higher-symmetry point groups, for monoclinic groups see the supplementary material (SM) Roig et al. .
The lowest order invariant coupling $M_{i}$ and $N_{j}$ can be determined from symmetry analysis. The magnetization ${\vec{M}}$ belongs to the axial vector irreducible representation (IR) $\Gamma_{A}$, see SM Roig et al. . In contrast, ${\vec{N}}$ belongs to the IR $\Gamma_{A}\otimes\Gamma_{N}$, with $\Gamma_{N}$ the IR denoting the symmetry of the AM spin splitting. Hence, a coupling of $M_{i}$ and $N_{j}$ exists if the direct product $\Gamma_{A}\otimes\Gamma_{A}\otimes\Gamma_{N}$ contains the IR transforming trivially under all point group operations or, equivalently, there is a free energy invariant if $\Gamma_{A}\otimes\Gamma_{A}$ contains $\Gamma_{N}$ McClarty and Rau (2024). In Table 1 we provide the form of the free energy invariants considering the relevant point groups and symmetries for the spin splitting identified in Ref. Roig et al. (2024). Note that for $C_{6h}$, $O_{h}$ and certain IRs $\Gamma_{N}$ of $D_{6h}$, a bilinear coupling is not allowed, and therefore the coupling is of higher order. Following a similar procedure, the lowest-order coupling between the magnetization and the Néel vector can also be identified Hecker et al. (2024); Roig et al. , and is included in Table 1. Hence, this implies that for MnTe and CrSb, which have point group $D_{6h}$ and $\Gamma_{N}=B_{1g}$, the FM moment will be small, in agreement with Refs. Mazin (2024); Autieri et al. (2023).
Table 1: Free energy invariant for the bilinear coupling between $M_{i}$ and $N_{j}$ in the presence of SOC, with $P$ denoting the point group and $\Gamma_{N}$ corresponding to the IR of the spin splitting in each point group. The last column specifies if the coupling is generated from the order parameter $\tau_{z}\sigma_{i}$, or a secondary order parameter is needed to obtain the coupling.
@@
Microscopic models and secondary order parameters.– To investigate the dependence of the induced FM spin moment on the SOC strength, we consider the general form of the minimal model for AM from Ref. Roig et al. (2024) with the normal state Hamiltonian
@@
where $\tau_{i}$ represent sublattice and $\sigma_{i}$ spin degrees of freedom. The specific form of the parameters entering the model depends on the space group, the point group and the Wyckoff site symmetry. The inter-sublattice hopping term $t_{x,{\bf k}}$ exhibits the full symmetry of the point group. However, the crystal asymmetric hopping term $t_{z,{\bf k}}$ must exhibit a ${\bf k}$ dependence that transforms as the non-trivial IR $\Gamma_{N}$ since it describes the local symmetry breaking from multipole moments  Hayami et al. (2018); Bhowal and Spaldin (2024); Roig et al. (2024). To study the interplay of the Néel order ${\vec{N}}$ and the induced ${\vec{M}}$, we consider the perturbation
@@
where $G_{0}^{(\pm)}({\bf k},i\omega_{n})=\frac{1}{i\omega_{n}-(\varepsilon_{0,{\bf k}}\pm\tilde{E}_{{\bf k}})}$ denotes the Green’s function in the band basis, with the two-fold degenerate eigenenergies ${E}_{\bf k}^{\pm}=\varepsilon_{0,{\bf k}}\pm\tilde{E}_{{\bf k}}$ with $\tilde{E}_{{\bf k}}=\sqrt{t_{x,{\bf k}}^{2}+t_{z,{\bf k}}^{2}+\vec{\lambda}_{{\bf k}}^{2}}$. The projection operator |$\ket{u^{a}_{\bf k}}\bra{u_{\bf k}^{a}}$ | in Eq. (5) transforms from sublattice basis onto band $a$ at wavevector ${\bf k}$ Graf and Piéchon (2021); Roig et al. .
To examine the bilinear coupling between ${\vec{N}}$ and ${\vec{M}}$, we derive an expression for the coefficient $c_{ij}$ in Eq. (1). The analytic expressions for the other coefficients in Eq. (1) can be found in the SM Roig et al. . The quadratic free energy contribution in Eq. (4) coupling ${\vec{M}}$ and ${\vec{N}}$ is given by
@@
incorporating the density of states and Lindhard term. The combination $({\vec{M}}\crossproduct{\vec{N}})$ in Eq. (7) reveals that a nonzero invariant exists only if the antisymmetric direct product of the two axial IRs $[\Gamma_{A}\otimes\Gamma_{A}]_{-}$ contains $\Gamma_{N}$. In Table 1 we list whether the invariant can be generated to linear order of SOC, the antisymmetric product for the different point groups in detailed in the SM Roig et al. . The SOC is expected to have a weak effect in AM Šmejkal et al. (2022a), and therefore the induced magnetization will be vanishingly small when it is not generated to linear order. As seen from Table 1, for the point group $D_{4h}$ the SOC-linear invariant is only generated for $\Gamma_{N}=A_{2g}$. Focusing on crystals with rutile structure i.e. space group (SG) 136 and Wyckoff position 2a for the magnetic atoms as in RuO2, MnF2, NiF2, and CoF2, we have $\Gamma_{N}=B_{2g}$. Consequently, the induced ${\vec{M}}$ is at least quadratic in SOC, as opposed to the material candidates Nb2FeB2 and Ta2FeB2, which have $\Gamma_{N}=A_{2g}$. Notably, Table 1 also shows that the FM moment induced in orthorhombic materials ($D_{2h}$) is generally expected to be larger.
@@
To further verify these points, in Fig. 1 we show the calculated ${\vec{M}}=\mu_{B}\sum_{a,{\bf k}}\bra{u^{a}_{\bf k}}\vec{S}\ket{u^{a}_{\bf k}}f(E^{a}_{{\bf k}})$ relevant for (a) rutile structure ($D_{4h}$, $\Gamma_{N}=B_{2g}$) and (b) FeSb2 structure ($D_{2h}$, $\Gamma_{N}=B_{1g}$). As seen, the SOC-induced ${\vec{M}}$ indeed scales quadratically (linearly) with the SOC strength for SG 136 (FeSb2) and is significantly smaller for the band relevant for rutile AM compared to FeSb2. For the case of FeSb2 we additionally compare the induced $\vec{M}$ for the Néel vector along the $x$-axis and the $y$-axis in Fig. 1(b) to see if $\vec{M}$ follows the predicted microscopic $(M_{x}N_{y}-M_{y}N_{x})$ result for the invariant. Indeed, even though $M_{x}$ and $M_{y}$ are not symmetry-related, they are of opposite sign and nearly identical in magnitude for this material. In summary, these results explicitly demonstrate the properties summarized in Table 1, which can be applied to gain similar insight for many classes of AM materials.
Finally, we note that in order to understand the quadratic dependence of $\vec{M}$ on SOC, when $\vec{M}$ is not allowed to linear order, requires the inclusion of secondary order parameters McClarty and Rau (2024). Specifically, when AM order sets in, secondary order parameters are also induced by symmetry, and they can give rise to a finite coupling between the two orders ${\vec{M}}$ and ${\vec{N}}$. Such secondary order parameters include other spin textures as well as a current loop order, see SM Roig et al. . The free energy for a secondary order parameter $\vec{O}$ can be written as
@@
which couples bilinearly to ${\vec{M}}$ and ${\vec{N}}$. Thus, to higher order ${\vec{N}}$ can also induce a magnetization ${\vec{M}}$ through the secondary order parameter $\vec{O}$. Our minimal model suggests that in this case the FM spin moment is at least quadratic in SOC.
Quasi-symmetry protection of negligible FM.– It is possible to understand the above results using the recently introduced concept of quasi-symmetry Li et al. (2024b); Guo et al. (2022), describing emergent approximate symmetries when certain terms in the Hamiltonian become negligible. Here we use SOC to generate our quasi-symmetry. Specifically, we consider the quasi-symmetry that emerges when two of the SOC components ($\lambda_{x},\lambda_{y},$ or $\lambda_{z}$) vanish, as illustrated in Fig. 2 for a tetragonal system. This is relevant for the Landau coefficients since the SOC-linear contributions can be expressed as a linear combination of $\lambda_{x,y,z}$. For example, we consider that the $\lambda_{z}$-linear contribution is non-zero when the other two SOC terms are neglected. In that case, the normal state Hamiltonian gains additional symmetries, i.e., quasi-symmetries. Importantly, whether the $\lambda_{z}$-linear contribution to the Landau coefficient is permitted depends not only on the intrinsic symmetries of the crystal, but also on these emergent quasi-symmetries Li et al. (2024b); Guo et al. (2022).
@@
Here, we apply this to the Landau coefficients of $M_{x}N_{y}$ and $M_{y}N_{x}$. When only $\lambda_{x}$ ($\lambda_{y}$) SOC is present, the normal state Hamiltonian acquires an additional two-fold spin rotational symmetry $C^{\rm spin}_{2x}$ ($C^{\rm spin}_{2y}$), see Fig. 2(a) (Fig. 2(b)), which prohibits $\lambda_{x}$ (or $\lambda_{y}$)-linear contribution to these two Landau coefficients. When only $\lambda_{z}$ is present, a relevant quasi-symmetry is 4-fold spin rotational symmetry $C^{\rm spin}_{4z}$, as seen from Fig. 2(c). Under this symmetry, the SOC-linear contribution to $M_{x}N_{y}$ and $M_{y}N_{x}$ coefficients must have opposite sign. This behavior is universal and applies to any space group, regardless of whether it has intrinsic 4-fold rotational symmetry, as demonstrated in Fig. 1(b) for FeSb2. In structures with SG 136, however, its intrinsic tetragonal crystal symmetries force these two coefficients to be identical at all orders of SOC. Consequently, the SOC-linear contribution must vanish, as shown in Fig. 1(a). The complete discussion on other space groups can be found in SM Roig et al. . Hence, the SOC enabled quasi-symmetry demonstrates that the microscopic model leads to a general result, and naturally explains when the FM spin moment is large or small depending on the AM symmetry. The only assumption on the microscopic Hamiltonian that underlies this analysis is that AM is an instability purely in the spin-channel, that is, it is driven only by exchange interactions without any orbital angular momentum contribution.
Anomalous Hall effect.– In the SM we provide a detailed analysis of the AHE. There are two main conclusions; the first is that it is generically nonzero when allowed by symmetry and the second is that it is dominated by the following term in the Berry curvature Roig et al. (2024),
$\Omega_{\alpha,\beta,ij}^{(N_{x})}=$ $\frac{1}{2E_{\alpha,\beta}^{3}}\sum_{m,n=i,j}\varepsilon_{mn}\Big{[}(N_{x}+\beta|t_{z,{\bf k}}|)\partial_{m}\lambda_{x,{\bf k}}\partial_{n}t_{x,{\bf k}}$
$+\text{sgn}(t_{z,{\bf k}})\beta\,t_{x,{\bf k}}\partial_{m}t_{z,{\bf k}}\partial_{n}\lambda_{x,{\bf k}}$
$+\text{sgn}(t_{z,{\bf k}})\beta\,\lambda_{x,{\bf k}}\partial_{m}t_{x,{\bf k}}\partial_{n}t_{z,{\bf k}}\Big{]},$ (10)

### v2

@@ Abstract
Altermagnets break time-reversal symmetry and their spin-orbit coupling (SOC) allow for an anomalous Hall effect (AHE) that depends on the direction of the Néel ordering vector. The AHE and the ferromagnetic spin moment share the same symmetry and hence are usually proportional. However, density functional theory (DFT) calculations find that the AHE exists with negligible ferromagnetic spin moment for some compounds, whereas it reaches sizable values for other altermagnets. By examining realistic minimal models for altermagnetism in which the DFT phenomenology is captured, we uncover a general SOC-enabled quasi-symmetry, the uniaxial spin space-group, that provides a natural explanation for the amplitude of the ferromagnetic spin moment across the vast range of different altermagnetic materials. Additionally, we derive analytic expressions for the magnetic anisotropy energy, providing a simple means to identify the preferred Néel vector orientation for altermagnets.
@@
Introduction.– The Hall effect has been a frontier theme in condensed matter physics for more than half a century. Fundamental understanding of the quantized Hall conductance and the anomalous Hall effect (AHE) has helped pave the way for topological classifications of quantum matter and the importance of Berry curvature in transport properties Nagaosa et al. (2010); Sodemann and Fu (2015); Anyuan Gao, Yu-Fei Liu, Jian-Xiang Qiu, Barun Ghosh, Thaís V. Trevisan, Yugo Onishi, Chaowei Hu, Tiema Qian, Hung-Ju Tien, Shao-Wen Chen, Mengqi Huang, Damien Bérubé, Houchen Li, Christian Tzschaschel, Thao Dinh, et al. (2023); Kaplan et al. (2024); Fang et al. (2024); Du et al. (2021); Wang et al. (2023). The discovery of altermagnetism (AM) and its associated large antiferromagnetic AHE is yet another testimony to this development Lee et al. (2024); Krempaský et al. (2024); Osumi et al. (2024); Li et al. (2024a); Yang et al. (2025); Reimers et al. (2024); Ding et al. (2024); Feng et al. (2022); Kluczyk et al. (2024); Gonzalez Betancourt et al. (2023), which may open new opportunities for devices utilizing dissipationless transport and spintronics technologies Šmejkal et al. (2020, 2022a, 2022b, 2022c); Fang et al. (2024); Ma et al. (2021).
@@
The AHE in AM, which we describe here as an intra-unit cell Néel order, is only nonzero for certain directions of the Néel vector and in the presence of relativistic spin-orbit coupling (SOC). This, in addition, induces weak ferromagnetism (FM), which satisfies the same symmetry properties as the AHE McClarty and Rau (2024); Xiao et al. (2024). This implies that, in principle, FM and AHE share the same dependence on the orientation of the Néel vector Šmejkal et al. (2022c); Fernandes et al. (2024); Cheong and Huang (2024). However, from DFT it is found that the relative amplitude of the AHE transport coefficient and the FM spin moment depends strongly on the particular AM material under investigation. For example, for RuO2 and MnTe, DFT finds a large AHE of 40 S/cm and 300 S/cm, respectively, while RuO2 has a nearly zero FM moment (that we estimate is of the order $10^{-7}\mu_{B}$) and Mn in MnTe has $10^{-4}\mu_{B}$mejkal et al. (2020); Mazin and Belashchenko (2024); Gonzalez Betancourt et al. (2023). In the case of MnTe and RuO2, weak FM was also reported experimentally Berlijn et al. (2017); Kluczyk et al. (2024), though the AM ground state of the latter material is still debated Hiraishi et al. (2024); Keßler et al. (2024). A weak FM spin moment has also been identified by DFT in CrSb Autieri et al. (2025). By contrast, in other AM, such as FeSb2, DFT finds a large AHE of 143 S/cm Mazin et al. (2021); Attias et al. (2024), and our calculations estimate a moderate FM moment of 0.03 $\mu_{B}$. Similarly, the material RuF4 has also been predicted by DFT to exhibit a large FM spin moment of 0.22 $\mu_{B}$ on the Ru sites Milivojević et al. (2024). These results suggest that while the AHE is generically large when allowed by symmetry, the corresponding FM moment is strongly dependent on the AM state and this is poorly understood.
@@
To gain deeper insight into the interplay between the AHE, the FM moment, and the AM state it is crucial to analyze the role of SOC using realistic models. Here, by examining realistic microscopic models for AM in which the DFT phenomenology is captured Roig et al. (2024), we identify which AM states have a large SOC induced AHE but a small FM spin moment and find good agreement with DFT results. In addition, we find that our model-based results are more generally valid. Specifically, they apply to all microscopic theories in which AM appears as a consequence of the exchange interaction. To show this, we identify a general SOC-enabled quasi-symmetry Li et al. (2024b); Guo et al. (2022), a uniaxial spin-space group, that establishes for which AM symmetries the symmetry-allowed FM is large or vanishingly small. Finally, the presence of SOC in AM breaks the spin-space degeneracy and leads to a preferred direction for the Néel vector. For example, DFT calculations have revealed that the moments are orthogonal to the a-b plane in RuO2 Feng et al. (2022), while in MnTe, FeSb2, Nb2FeB2 and Ta2FeB2 the moments are predicted to be in-plane Mazin et al. (2021); Mazin and Belashchenko (2024); Hou et al. (2023). By examining analytic expressions for the Landau coefficients derived from realistic microscopic models, we elucidate how the structure of SOC allows for the AM anisotropy energy to be understood.
@@
Interplay between magnetization and AM Néel order.– Introducing ${\vec{N}}$ as the intra-unit cell Néel order and ${\vec{M}}$ as the magnetization, the general form of the free energy density in the presence of SOC can be written as
@@
where $a_{N}$, $b_{N}$, $a_{M}$ and $c_{ij}$ are temperature-dependent Landau coefficients, $\vec{h}$ is an applied field, and $s_{1}$ and $s_{2}$ are the coefficients determining the magnetic anisotropy energy due to SOC. Note that the bilinear coupling between the two orders $M_{i}$ and $N_{j}$ is only allowed in the presence of SOC McClarty and Rau (2024). Without SOC, the spin-space group rotations, $[R_{S}||R_{G}]$, with spin space ($R_{S}$) and real space group operations ($R_{G}$) are uncoupled, which forces the bilinear coupling to vanish since $\vec{N}$ is non-trivial under pure $R_{G}$ operations. In the magnetic space group, the rotations act simultaneously on both spaces, and therefore $R_{S}$ and the rotation portion of $R_{G}$ must be the same, allowing this bilinear coupling to be non-zero. Eq. (1) describes orthorhombic or higher-symmetry point groups, for monoclinic groups see the supplementary material (SM) Roig et al. .
Table 1: Lowest order free energy invariant between $M_{i}$ and $N_{j}$ in the presence of SOC for different point groups $P$ and IRs of the spin splitting $\Gamma_{N}$, with the function $f_{\Gamma_{N}}({\bf k})$ transforming as $\Gamma_{N}$. The last column indicates if the $M_{i}$ generated from the AM order parameter is linear order in the SOC, denoted by a check mark, or higher order in SOC, denoted by a cross.
@@
The lowest order invariant coupling $M_{i}$ and $N_{j}$ can be determined from symmetry analysis. The magnetization ${\vec{M}}$ belongs to the axial vector irreducible representation (IR) $\Gamma_{A}$, see SM Roig et al. . In contrast, ${\vec{N}}$ belongs to the IR $\Gamma_{A}\otimes\Gamma_{N}$, with $\Gamma_{N}$ the IR denoting the symmetry of the AM spin splitting. Hence, a coupling of $M_{i}$ and $N_{j}$ exists if the direct product $\Gamma_{A}\otimes\Gamma_{A}\otimes\Gamma_{N}$ contains the IR transforming trivially under all point group operations or, equivalently, there is a free energy invariant if $\Gamma_{A}\otimes\Gamma_{A}$ contains $\Gamma_{N}$ McClarty and Rau (2024). In Table 1 we provide the form of the free energy invariants considering the relevant point groups and symmetries for the spin splitting identified in Ref. Roig et al. (2024). Note that for $C_{6h}$, $O_{h}$ and certain IRs $\Gamma_{N}$ of $D_{6h}$, a bilinear coupling is not allowed, and therefore the coupling is of higher order. Following a similar procedure, the lowest-order coupling between the magnetization and the Néel vector can also be identified Hecker et al. (2024); Roig et al. , and is included in Table 1.
Microscopic models.– To investigate the dependence of the induced FM spin moment on the SOC strength, we initially consider the general form of the minimal model for AM from Ref. Roig et al. (2024), and carry out a model-independent quasi-symmetry based analysis later. Thus, we start from the normal state Hamiltonian
@@
where $\tau_{i}$ represent sublattice and $\sigma_{i}$ spin degrees of freedom. The specific form of the parameters entering the model depends on the space group, the point group and the Wyckoff site symmetry. Here, $t_{x,{\bf k}}$ is an inter-sublattice hopping term, $\varepsilon_{0,{\bf k}}$ is the sublattice independent dispersion, and $\vec{\lambda}_{\bf k}$ is the SOC. The crystal asymmetric hopping term $t_{z,{\bf k}}$ exhibits a ${\bf k}$ dependence that transforms as the non-trivial IR $\Gamma_{N}$ since it describes the local symmetry breaking from multipole moments  Hayami et al. (2018); Bhowal and Spaldin (2024); Roig et al. (2024).
Initially, we analyze the AHE. For the Néel order along the an arbitrary direction $l$, $N_{l}$, the only contribution to the Berry curvature that is linear in SOC originates from the SOC component parallel to $l$, that is $\lambda_{l}$, and is given by Roig et al. (2024, )
$\Omega_{\alpha,\beta,ij}^{(N_{l})}=$ $\frac{1}{2E_{\alpha,\beta}^{3}}\sum_{m,n=i,j}\varepsilon_{mn}\Big{[}(N_{l}+\beta|t_{z,{\bf k}}|)\partial_{m}\lambda_{l,{\bf k}}\partial_{n}t_{x,{\bf k}}$
$+\text{sgn}(t_{z,{\bf k}})\beta\,t_{x,{\bf k}}\partial_{m}t_{z,{\bf k}}\partial_{n}\lambda_{l,{\bf k}}$
$+\text{sgn}(t_{z,{\bf k}})\beta\,\lambda_{l,{\bf k}}\partial_{m}t_{x,{\bf k}}\partial_{n}t_{z,{\bf k}}\Big{]},$ (3)
where the dispersion is $E_{\alpha=\pm,\beta=\pm}=\alpha\Big{(}N_{l}^{2}+\lambda_{l,{\bf k}}^{2}+t_{x,{\bf k}}^{2}+t_{z,{\bf k}}^{2}+2\beta|N_{l}||t_{z,{\bf k}}|\Big{)}^{1/2}$. This expression is generically non-zero. Thus, when allowed by symmetry, the AHE is expected to be large and linear in SOC for all point groups and AM symmetries. In the SM Roig et al. , we carry out calculations of the AHE that support this conclusion.
To study now the interplay of the Néel order ${\vec{N}}$ and the induced ${\vec{M}}$, we consider the perturbation
@@
where $G_{0}^{(\pm)}({\bf k},i\omega_{n})=\frac{1}{i\omega_{n}-(\varepsilon_{0,{\bf k}}\pm\tilde{E}_{{\bf k}})}$ denotes the Green’s function in the band basis, with the two-fold degenerate eigenenergies ${E}_{\bf k}^{\pm}=\varepsilon_{0,{\bf k}}\pm\tilde{E}_{{\bf k}}$ with $\tilde{E}_{{\bf k}}=\sqrt{t_{x,{\bf k}}^{2}+t_{z,{\bf k}}^{2}+\vec{\lambda}_{{\bf k}}^{2}}$. The projection operator $P_{\bf k}^{a}=\ket{u^{a}_{\bf k}}\bra{u_{\bf k}^{a}}$ | in Eq. (6) transforms from the sublattice basis onto band $a$ at wavevector ${\bf k}$ Graf and Piéchon (2021); Roig et al. .
To examine the bilinear coupling between ${\vec{N}}$ and ${\vec{M}}$, we derive an expression for the coefficient $c_{ij}$ in Eq. (1). The analytic expressions for the other coefficients in Eq. (1) can be found in the SM Roig et al. . At one-loop level, the quadratic free energy contribution in Eq. (5) coupling ${\vec{M}}$ and ${\vec{N}}$ is given by
@@
incorporating the density of states and Lindhard term. The combination $({\vec{M}}\crossproduct{\vec{N}})$ in Eq. (8) reveals that a nonzero invariant exists only if the antisymmetric direct product of the two axial IRs $[\Gamma_{A}\otimes\Gamma_{A}]_{-}$ contains $\Gamma_{N}$. In Table 1 we list whether the invariant can be generated to linear order of SOC, the antisymmetric product for the different point groups is detailed in the SM Roig et al. . The SOC is expected to be a weak effect in AM Šmejkal et al. (2022a), and therefore the induced magnetization will be vanishingly small when it is not generated to linear order. As seen from Table 1, for the point group $D_{4h}$ the SOC-linear invariant is only generated for $\Gamma_{N}=A_{2g}$. Focusing on crystals with rutile structure i.e. space group (SG) 136 and Wyckoff position 2a for the magnetic atoms as in RuO2, MnF2, NiF2, and CoF2, we have $\Gamma_{N}=B_{2g}$. Consequently, the induced ${\vec{M}}$ is at least quadratic in SOC, as opposed to the material candidates Nb2FeB2 and Ta2FeB2, which have $\Gamma_{N}=A_{2g}$. Notably, Table 1 also shows that the FM moment induced in orthorhombic materials ($D_{2h}$) is generally expected to be larger.
@@
To further verify these points, in Fig. 1 we show the calculated ${\vec{M}}=\mu_{B}\sum_{a,{\bf k}}\bra{u^{a}_{\bf k}}\vec{S}\ket{u^{a}_{\bf k}}f(E^{a}_{{\bf k}})$ relevant for (a) rutile structure ($D_{4h}$, $\Gamma_{N}=B_{2g}$) and (b) FeSb2 structure ($D_{2h}$, $\Gamma_{N}=B_{1g}$). As seen, the SOC-induced ${\vec{M}}$ indeed scales quadratically (linearly) with the SOC strength for SG 136 (FeSb2) and is significantly smaller for the band relevant for rutile AM compared to FeSb2. For the case of FeSb2 we additionally compare the induced $\vec{M}$ for the Néel vector along the $x$-axis and the $y$-axis in Fig. 1(d) to see if $\vec{M}$ follows the predicted microscopic $(M_{x}N_{y}-M_{y}N_{x})$ result for the invariant. Indeed, even though $M_{x}$ and $M_{y}$ are not symmetry-related in this orthorhombic system, they are of opposite sign and nearly identical in magnitude for this material. In summary, these results explicitly demonstrate the properties summarized in Table 1, which can be applied to gain similar insight for many classes of AM materials.
Finally, we note that in order to understand the quadratic dependence of $\vec{M}$ on SOC from exact diagonalization, when $\vec{M}$ is not allowed to linear order, requires the inclusion of secondary order parameters McClarty and Rau (2024), equivalent to two-loop calculations  Roig et al. . Specifically, when AM order sets in, secondary order parameters are also induced by symmetry, and they can give rise to a finite coupling between the two orders ${\vec{M}}$ and ${\vec{N}}$. Such secondary order parameters include other spin textures as well as a current loop order, see SM Roig et al. . The free energy for a secondary order parameter $\vec{O}$ can be written as
@@
which couples bilinearly to ${\vec{M}}$ and ${\vec{N}}$. Thus, ${\vec{N}}$ can also induce a magnetization ${\vec{M}}$ through the secondary order parameter $\vec{O}$, and our minimal models reveal that in this case the FM spin moment is at least quadratic in SOC, as discussed in the SM Roig et al. .
Quasi-symmetry protection of negligible FM.– Our minimal models are in agreement with the DFT results, which suggests a more general explanation beyond specific loop expansions or microscopic models. Indeed, it is possible to understand the above results using the recently introduced concept of quasi-symmetry Li et al. (2024b); Guo et al. (2022), describing emergent approximate symmetries when certain terms in the Hamiltonian become negligible. Here we use SOC to generate our quasi-symmetry. Specifically, we consider the quasi-symmetry that emerges when two of the SOC components ($\lambda_{x},\lambda_{y},$ or $\lambda_{z}$) vanish, as illustrated in Fig. 2 for a tetragonal system. The resulting symmetry group, which we denote the uniaxial spin-space group due to its spin rotational invariance around the SOC direction, has higher symmetry than the magnetic space group but lower symmetry than the spin space group. This is relevant for determining which Landau coefficients are linear in one of $\lambda_{x,y,z}$. For example, for $\lambda_{z}\neq 0$ with $\lambda_{x}=\lambda_{y}=0$, the normal state Hamiltonian gains additional symmetries, i.e., quasi-symmetries. Importantly, whether the $\lambda_{z}$-linear contribution to the Landau coefficient is permitted depends not only on the intrinsic symmetries of the crystal, but also on these emergent quasi-symmetries Li et al. (2024b); Guo et al. (2022).
@@
Here, we apply this to the Landau coefficients of $M_{x}N_{y}$ and $M_{y}N_{x}$. When only $\lambda_{x}$ ($\lambda_{y}$) SOC is present, the normal state Hamiltonian acquires an additional two-fold spin-rotational symmetry $[C_{2x}||E]$ ($[C_{2y}||E]$), see Fig. 2(a) (Fig. 2(b)), which prohibits $\lambda_{x}$ (or $\lambda_{y}$)-linear contribution to these two Landau coefficients since $M_{x}$ ($N_{x}$) is odd under this quasi-symmetry while $N_{y}$ ($M_{y}$) is even. When only $\lambda_{z}$ is present, a relevant quasi-symmetry in the uniaxial spin point group is the four-fold spin-rotational symmetry $[C_{4z}||E]$. As seen from Fig. 2(c), under this symmetry the SOC-linear contribution to $M_{x}N_{y}$ and $M_{y}N_{x}$ coefficients must have opposite sign. This behavior is universal and applies to any space group, regardless of whether it has intrinsic 4-fold rotational symmetry, as demonstrated in Fig. 1(d) for FeSb2. In structures with SG 136, however, its intrinsic tetragonal crystal symmetries force these two coefficients to be identical at all orders of SOC. Consequently, the SOC-linear contribution must vanish, as shown in Fig. 1(b). The complete discussion on other space groups can be found in SM Roig et al. . This also implies that for MnTe and CrSb, which have point group $D_{6h}$ and $\Gamma_{N}=B_{1g}$, the FM moment will be cubic in SOC (see End Matter), independent of the non-linear coupling terms in Table 1, in agreement with Refs. Mazin and Belashchenko (2024); Autieri et al. (2025). Hence, the SOC enabled quasi-symmetry demonstrates that the results of the microscopic model are very general, and naturally explains when the FM spin moment is large or small depending on the AM symmetry. The only assumption on the microscopic Hamiltonian that underlies this analysis is that AM is an instability purely in the spin-channel, that is, it is driven only by exchange interactions without any orbital angular momentum contribution.